1/15/15 5:28pm update RV

The issue has been resolved. The replacement switch has been installed and configured successfully.

Thank You,

ITS Help Desk


1/15/15 1:58pm update DR

A confirmed electrical outage this morning caused the problem. One switch is dead and all efforts to restore it was futile. A replacement switch is being installed and configured.

ETA – approximately 2pm service should be restored.


Thank you,

ITS Help Desk


1/15/15 9:44am update DR

Many devices went down and/or rebooted at the DOSA site (1510 Waters Place) this morning.

Connectivity to the server switch still appears down. Networking is currently investigating, and there is no ETA.

Further updates will be posted as they become available.

Thank you,

ITS Help Desk


Comments are closed.